OpenStudy (anonymous):

12. What is the 7th term of the geometric sequence where a1 = -4 and a5 = -1,024?

OpenStudy (mathstudent55):

The n-th term of a geometric sequence is: \[a _{n} = a _{1} r ^{n-1}\] You have \[a _{5} = a _{1} r ^{5-1}\] -1024 = -4(r^4) r^4 = 256 r = 4 Now that you know r, use the formula for the n-th term with a1=-4 and n = 7.

OpenStudy (anonymous):

-21844?

OpenStudy (mathstudent55):

a7 = a1(r^6) a7 = -4(4^6) a7 = -16384